# MDMP: Multi-modal Diffusion for supervised Motion Predictions with uncertainty

### v1

@@ Abstract
This paper introduces a Multi-modal Diffusion model for Motion Prediction (MDMP) that integrates and synchronizes skeletal data and textual descriptions of actions to generate refined long-term motion predictions with quantifiable uncertainty. Existing methods for motion forecasting or motion generation rely solely on either prior motions or text prompts, facing limitations with precision or control, particularly over extended durations. The multi-modal nature of our approach enhances the contextual understanding of human motion, while our graph-based transformer framework effectively capture both spatial and temporal motion dynamics. As a result, our model consistently outperforms existing generative techniques in accurately predicting long-term motions. Additionally, by leveraging diffusion models’ ability to capture different modes of prediction, we estimate uncertainty, significantly improving spatial awareness in human-robot interactions by incorporating zones of presence with varying confidence levels for each body joint.
@@ 1 Introduction
Through collaboration and assistance, robots could significantly augment human capabilities across diverse sectors, including smart manufacturing, healthcare, agriculture, and construction. Indeed, they can complement the critical and adaptive decision-making skills of human workers with higher precision and consistency in repetitive tasks. However, one challenge prohibiting human-robot collaboration is the safety of workers in the presence of robots. To act safely and effectively together, continuous knowledge of future human motion and location in the common workspace with a measure of uncertainty is pivotal. This real-time awareness allows robots to adjust their trajectories to avoid collision and perform precise collaborative tasks [25, 58, 5].
Figure 1: Application in Human-Robot Collaboration: Based on the action: ”Picking up an object; giving object to the robot” our MDMP predicts the future movements of the human worker with an uncertainty zone of presence around each body joint.
Figure 2: SMPL Meshes of MDMP Predicted motions of different scenarios. The text descriptions vertically associated to the motions as well as the blue frames are the inputs of the model. The orange frames are the predictions, darker colors indicate later frames.
Humans can predict future events based on their self-constructed models of physical and socio-cultural systems. This skill, developed from childhood through observation and active participation in society, enables them to anticipate others’ movements. Researchers are now trying to transfer this capability often refered as ”Theory of mind” [10] to machines by training them to learn similar motion estimation tasks. Current methodologies fall into two primary categories: Human Motion Forecasting (see Section 2.2) and Human Motion Generation (see Section 2.3). While the former uses only a short input sequence of skeletal motion to predict its future trajectory, the latter relies exclusively on textual prompts to generate motion sequences. Despite advancements in text-to-motion models, challenges remain in controlling generation due to the expansive action space a simple prompt can describe, which may not always align with human expectations or behavior.
Conversely, motion prediction using past sequences is a long-standing challenge that has achieved high accuracy over short-term predictions but struggles with long-term predictions. This is understandable, as even for humans, predicting someone’s immediate future movement based on past motions is feasible, but beyond one or two seconds, the variety of possibilities makes it nearly impossible without context. However, knowledge of the intended action provides a rough idea of future positions, as the contextual information of the action guides intuitions.
@@
In this work, inspired by MDM [50] (Motion Diffusion Model) and LTD [36] (Learning Trajectory Dependencies), we propose a transformer-based diffusion model with a Graph Convolutional Encoder optimized for the spatio-temporal dynamics of motion data. A key design element is the use of learnable graph connectivity, as introduced by Mao et al. [36], to more effectively capture joint dependencies. Our model is trained in a classifier-free manner [18], allowing for a balance between variability and accuracy, with an emphasis on precision to ensure safety in HRC. Additionally, our Multi-modal Diffusion Model for Motion Predictions (MDMP) harnesses the stochastic nature of diffusion models to predict presence zones with varying confidence levels. This uncertainty measure is particularly crucial for long-term motion predictions, where uncertainty grows over time. By offering a spatial understanding of human presence, our model significantly enhances collision avoidance, improving safety and real-time interaction in dynamic collaboration scenarios.
We summarize the contributions as follows: 1) A novel multi-modal diffusion model integrating textual and skeletal data to generate long-term motion predictions. 2) A graph-based transformer framework that effectively captures spatial and temporal dynamics of motion. 3) A method for estimating uncertainty in predictions, improving spatial awareness and safety in HRC scenarios. 4) A thorough validation of the predicted uncertainty through a comparative study and finally our implementation is open-sourced.
@@ 2.3 Human Motion Generation
Instead of predicting future motion based on past sequences, some generative methods are conditioned natural language [1, 41] to overcome this short-term issue. This approach faces other challenges such as the vast variability of possible motions corresponding to the same label. However, Text2Motion has garnered significant interest and varied successful approaches. TEMOS [41] and T2M [14] employ a VAE to map text prompts to a latent space distribution of language and motion. MotionGPT [21] furthers this by proposing a unified motion-language framework. Recently, MDM [50] leveraged recent advancements in DDPMs, proving that diffusion models are a better candidate for human motion generation, as they can retain the formation of the original motion sequence and thus allows them to easily apply more constraints during the denoising process. LDM [6] took it forward by performing Diffusion in the latent space and MoMask [13] & MMM [42] leverages Masked Transformers.
By fixing some parts of a motion sequence and filling in the gaps, some of these Text2Motion baselines such as MDM [50] and MotionGPT [21] propose a form of ”motion editing” by forcing their models to generate motions with preserved original data. Unlike these methods, which only edit motions during sampling, our approach trains the model with both textual prompts and motion sequence conditioning to learn contextuality and guide generation towards precise predictions. While these models are compared on diversity and multi-modality metrics, our goal is to minimize the distance between predictions and ground-truth for accurate predictions in HRC.
@@ 3 Methodology
We now explain the architecture of our proposed MDMP in detail. For an overview, please refer to Figure 3. Our method uses the building blocks of MDM [50], but with three main differences: (1) a denoising model that includes variance learning to increase log-likelihood and perform uncertainty estimates, (2) a GCN encoder with learnable graph connectivity, and (3) a learning framework that incorporates contextuality by synchronizing skeletal inputs with initial textual inputs.
@@ 3.1 Problem Formulation
A motion sample can be represented by a temporal skeleton sequence $X=\{p^{i}\}_{i=1}^{N}$ where a frame $p_{i}$ denotes a pose that can be modeled using different joint feature representations depending on the dataset (see Section 4.1). The simplest form that any representation can easily revert to without any loss of information is the joints’ position in 3D space where $p_{i}=\{x(1)_{i},...,x(J)_{i}\}$ with joints $x(j)\in\mathbb{R}^{M}$ where $M=3$ and $J$ being the total number of joints. Some parameterizations use rotation matrices ($M=9$), angle-axis ($M=4$), or quaternion ($M=4$) to represent each joint, while others include information such as angular and/or linear velocity.
Our goal is to generate a motion sample conditioned on both a prefix of that sequence $Y=\{p^{i}\}_{i=1}^{n}$ with $n<N$ and a textual prompt that will be encoded into a textual embedding $c$ by CLIP [44].
@@ 3.2 The Variational Diffusion Process
Then, during both training and inference, we use MDMP (see Fig. 2a) as the decoder—conditioned at each step by the previously mentioned inputs $Y$ and $c$—to progressively denoise $X_{t}$ from a standard Gaussian. Instead of predicting the noise $\epsilon_{0}$ as formulated in DDPM [17], we follow [45] and  [50] and predict the signal itself along with its variance: $\hat{X}_{0},V_{0}=\text{MDMP}(X_{t},t,Y,c)$
@@
One subtlety is that $L_{\text{simple}}$ provides no learning signal for variances, as Ho et al. [17] chose to fix the variance rather than learn it. However, in our specific application, we leverage learned variances to generate presence zones with varying confidence levels to help ensure safety in HRC scenarios.
@@ 3.3 Learning the Variances of the Denoising process
To learn the reverse process variances, our model outputs a vector $V_{0}$ of the same shape as $\hat{X}_{0}$, and—following Nichol et al. [39]—we parameterize the variance as an interpolation between $\beta_{t}$ and $\tilde{\beta}_{t}$ in the log domain by turning this output $V_{0}$ into $\Sigma_{\theta,t}$ as follows:
@@
3.4 Encoding the joint features
To encode the spatial structure of the pose features, we leverage GCNs [52]. Instead of relying on a predefined sparse graph, we follow Mao et al. [36] and learn the graph connectivity during training, thus essentially learning the dependencies between the different joint trajectories. To this end, we model the human body through the joint features as a fully-connected graph with $K$ nodes. The strength of the edges in this graph can then be represented by a weighted adjacency matrix $A\in\mathbb{R}^{K\times K}$. The graph convolutional encoder/decoder then takes as input a matrix $H^{(\text{in})}\in\mathbb{R}^{K\times F}$, with $F$ the number of features. In our case, the encoder takes as input the $K\times N$ matrix of joint features for each frame $n\in[0,N]$ of the sequence. Given this information and a set of trainable weights $W\in\mathbb{R}^{F\times\hat{F}}$, a graph convolutional layer outputs a matrix of the form: $H^{(\text{out})}=AH^{(\text{in})}W$. Where $A$ is the trainable weighted adjacency matrix. All operations are differentiable with respect to both $A$ and $W$, which allows training on both the weights and the joint feature connectivities.
@@ 3.5 Predicting Uncertainty
  • Predicted Variance: The final approach leverages the variance of the predicted distribution of each motion sequence $\Sigma_{\theta,0}$ learned during training.
Both the second and third methods produce outputs in the same dimensions as the model, including predictions for root height, root angular and linear velocity, as well as joint positions and velocities in the local reference of the root. To calculate a single uncertainty index for each joint at each timestep, we average all features associated to the same joint.
3.6 Model Specifics
As shown in Figure 3, upstream of the encoder, the motion sample $X_{t}=\{p^{i}_{t}\}_{i=1}^{N}$ is conditioned by the input motion $Y=\{p^{i}\}_{i=1}^{n}$ through masking. The noise time-step $t$ and the textual embedding $c$ are projected to the same dimensional latent space by separate feed-forward networks. These features, summed with a sinusoidal positional embedding, are fed to the transformer encoder-only backbone [51]. The backbone output is projected back to the original motion dimensions and serves as the prediction. When learning variances, joint features are encoded in a latent space with twice the dimensions to decode $\hat{x}_{0}$ and their associated variance $V_{0}$ separately. Our model is trained both conditionally and unconditionally on text, by randomly masking 10% of the text embeddings. This allows to balance diversity and text-fidelity during sampling (see Figure 3 (Right)).
@@ 4 Experiments and Results
In this section, we present the experimental setup and evaluation of our proposed model. We describe the dataset used for training and testing, outline the metrics for accuracy and uncertainty, and provide details on our model’s implementation. Finally, we present a comprehensive quantitative and qualitative evaluation, including comparisons with Text2Motion baselines that propose Motion-Editing, analysis of uncertainty parameters, and an ablation study to assess the effects of motion-text fusion and architectural design choices.
@@ 4.1 Dataset
To train and evaluate our model, we use the HumanML3D [14] dataset, which is the largest and most diverse collection of scripted human motions. It combines motion sequences from the HumanAct12 [15] and AMASS [35] datasets, processed to standardize the motions to 20 FPS with a maximum length of 10 seconds per sequence. HumanML3D comprises 14,616 motions with 44,970 descriptions, covering 5,371 distinct words, totaling 28.59 hours of motion data with an average length of 7.1 seconds and three textual descriptions per sequence. The dataset is split for training and evaluation. For evaluation, we filter the set to include only motions longer than 3 seconds, allowing us to condition the models on 2.5 seconds of motion and predict at least 0.5 seconds into the future. After filtering, the evaluation set contains 4,328 out of the initial 4,646 motion sequences.
Figure 4: (Left) Temporal evolution of error in predictions. Quantitative Resutls on HumanML3D over MPJPE [mm]. (Right) 3D Plots of Motion Predictions (orange) vs Ground truth (blue). Motion Sequence example associated to textual prompt:“from a standing position, the person slowly walks in circle, clockwise, then stops”. Paler shades represents earlier frames.
@@ 4.2 Metrics for Accuracy and Uncertainty
Figure 5: (Left) Sparsification Error Plot. Quantitative Results of the uncertainty parameters: The Mode Divergence index closely follows the Oracle curve, indicating the strongest alignment between uncertainty estimates and true errors. (Right) Joint Position Evolution over the Denoising Process. The position is progressively denoised until it converges to its final prediction. The fluctuations are used as a parameter for uncertainty.
@@ 4.3 Implementation Details
Our models were trained on an NVIDIA Titan V GPU over 1.7 days and on NVIDIA Tesla V100 GPU over 1.2 days with a batch size of 64. We used 8 layers of the Transformer Encoder with 4 multi-head attention for each, separated by a GeLU activation function and a dropout value of $0.1$. The joint features from $X_{t}$ were encoded in a latent dimension of 1024 when learning variances and 512 without learning variances. 1024 corresponds to the concatenation of the joint features of $\hat{X}_{0}$ [512] and $V_{0}$ [512]. To encode the text, we use a frozen CLIP-ViT-B/32 model. Each model was trained for 600K steps, after which a checkpoint was chosen that minimizes the MPJPE metric to be reported. Our generative process is conditioned by a motion input sequence of 50 frames which represents 2.5 seconds at 20FPS. We also set $\lambda=0.001$ to prevent $L_{\text{VLB}}$ from overwhelming $L_{\text{simple}}$. We evaluate our models with guidance-scale $\mu=2.5$ but as discussed in the Motion & Text ablation study Section 4.4 this can be adapted for specific applications (eg. short/long-term predictions).
To evaluate the effectiveness of our multimodal fusion approach, we compare against Motion Editing baselines MotionGPT [21] and MDM [50], both of which also utilize language and motion data and are trained on HumanML3D [14]. We implement their pretrained versions (open-sourced) and compare on the entire test set of HumanML3D using MPJPE. To perform predictions we provide each compared model with the same motion prequel sequence (50 frames) and compare the rest of the predicted sequence to the ground truth.
4.4 Quantitative & Qualitative Evaluation
Model Accuracy Evaluation over MPJPE: Unlike the implemented baselines MotionGPT [21] and MDM [50] that treat motion data as a masked input during sampling, our model is trained to leverage it as an additional supervision signal, which we find to lead to significantly enhances performance, especially over longer sequences. In Fig. 4 the temporal evolution chart shows that our model substantially outperforms these baselines in accuracy, with consistently lower MPJPE values over time and a more gradual increase in error compared to the sharp rise seen in MDM [50] and MotionGPT [21]. These results are also demonstrated qualitatively in the 3D plots Fig. 4 (Right) (see Appendix & Video for more examples) where our predictions align more closely with the ground truth, especially towards the end of the sequence. Indeed, both baselines’ outputs fail to follow the indicated trajectory (projection of the root joint in the XZ-plane) whereas our model follows the ”circle”, almost aligning with the ground truth on the last frame.

### v2

@@ Abstract
This paper introduces a Multi-modal Diffusion model for Motion Prediction (MDMP) that integrates and synchronizes skeletal data and textual descriptions of actions to generate refined long-term motion predictions with quantifiable uncertainty. Existing methods for motion forecasting or motion generation rely solely on either prior motions or text prompts, facing limitations with precision or control, particularly over extended durations. The multi-modal nature of our approach enhances the contextual understanding of human motion, while our graph-based transformer framework effectively capture both spatial and temporal motion dynamics. As a result, our model consistently outperforms existing generative techniques in accurately predicting long-term motions. Additionally, by leveraging diffusion models’ ability to capture different modes of prediction, we estimate uncertainty, significantly improving spatial awareness in human-robot interactions by incorporating zones of presence with varying confidence levels. Code: github.com/leob03/mdmp
@@ 1 Introduction
Through collaboration and assistance, robots could significantly augment human capabilities across diverse sectors, including smart manufacturing, healthcare, agriculture, construction and many others. Indeed, they can complement the critical and adaptive decision-making skills of human workers with higher precision and consistency in repetitive tasks. However, one challenge prohibiting human-robot collaboration is the safety of workers in the presence of robots. To act safely and effectively together, continuous knowledge of future human motion and location in the common workspace with a measure of uncertainty is pivotal. This real-time awareness allows robots to adjust their trajectories to avoid collision and perform precise collaborative tasks [25, 57, 5].
Figure 1: MDMP integrates skeletal motion and text to generate long-term motion predictions with uncertainty zones, shown in both skeletal and 3D human mesh formats.
Figure 2: SMPL [34] Meshes of MDMP Predicted motions of different scenarios. The text descriptions vertically associated to the motions as well as the blue frames are the inputs of the model. The orange frames are the predictions, darker colors indicate later frames.
Humans can predict future events based on their self-constructed models of physical and socio-cultural systems. This skill, developed from childhood through observation and active participation in society, enables them to anticipate others’ movements. Researchers are now trying to transfer this capability to machines by training them to learn similar motion estimation tasks. Current methodologies fall into two main categories: Human Motion Forecasting (Section 2.2) and Human Motion Generation (Section 2.3). While the former uses only a short input sequence of skeletal motion to predict its future trajectory, the latter relies exclusively on textual prompts to generate motion sequences.
Despite advancements in text-to-motion models, challenges remain in controlling generation due to the expansive action space a simple prompt can describe, which may not always align with human expectations or behavior. Moreover, while some text-to-motion methods have been adapted to perform tasks like motion editing or motion prediction by conditioning their generative process on motion data during sampling, our study demonstrates that, since they are only fed with textual prompts during training, our method consistently outperforms them in terms of accuracy metrics.
Conversely, motion prediction using past sequences is a long-standing challenge that has achieved high accuracy over short-term predictions but struggles with long-term predictions. Even for humans, predicting someone’s immediate future movement based on past motions is feasible, but beyond one or two seconds, the multitude of possibilities makes it nearly impossible without context. However, knowledge of the intended action provides a rough idea of future positions, as the contextual information of the action guides intuition.
@@
In this work, inspired by MDM [50] (Motion Diffusion Model) and LTD [37] (Learning Trajectory Dependencies), we propose a transformer-based diffusion model with a Graph Convolutional Encoder optimized for the spatio-temporal dynamics of motion data. A key design element is the use of learnable graph connectivity, as introduced by Mao et al. [37], to more effectively capture joint dependencies. Additionally, our Multi-modal Diffusion Model for Motion Predictions (MDMP) harnesses the stochastic nature of diffusion models to predict presence zones—bounded regions in 3D space that are guaranteed to contain the true position of a particular body joint—with varying size based on uncertainty. This measure of confidence is particularly crucial for long-term motion predictions, where uncertainty grows over time. By offering a spatial understanding of human presence, our model should significantly enhances collision avoidance, improving safety and real-time interaction in dynamic collaboration scenarios.
We summarize the contributions as follows: 1) A novel multi-modal diffusion model trained on both textual and skeletal data for precise long-term motion predictions. 2) An uncertainty estimation method to significantly enhance spatial awareness and safety in HRC scenarios. 3) A graph-based transformer capturing spatial-temporal dynamics effectively. 4) A comprehensive validation of uncertainty estimation, with an open-source implementation.
@@ 2.3 Human Motion Generation
Instead of predicting future motion based on past sequences, some generative methods are conditioned on natural language [1, 42] to overcome this short-term issue. This approach faces other challenges such as the vast variability of possible motions corresponding to the same label. However, Text2Motion has garnered significant interest and varied successful approaches. TEMOS [42] and T2M [14] employ a VAE to map text prompts to a latent space distribution of language and motion. MotionGPT [20] furthers this by proposing a unified motion-language framework. MDM [50] proved that diffusion models are a better candidate for human motion generation, as they can retain the formation of the original motion sequence and thus allows them to easily apply more constraints during the denoising process. Then, LDM [6] performed the Diffusion in the latent space and MoMask [15] leveraged Masked Transformers.
By fixing some parts of a motion sequence and filling in the gaps, some of these Text2Motion baselines such as MDM [50], MotionGPT [20] and MoMask [15] propose a form of ”motion editing” by forcing their models to generate motions with preserved original data. Unlike these methods, which only edit motions during sampling, our approach trains the model with both textual prompts and motion sequence conditioning to learn contextuality and guide generation towards precise predictions. While these models are compared on diversity and multi-modality metrics, our goal is to minimize the distance between predictions and ground-truth for accurate predictions in HRC.
@@ 3 Methodology
We now explain the architecture of our proposed MDMP in detail. For an overview, please refer to Figure 3. As part of the Diffusion Process MDMP progressively denoises a motion sample conditioned by the input motion through masking. Our architecture employs a GCN encoder to capture spatial joint features. We encode text prompts using CLIP followed by a linear layer; the textual embedding $c$ and the noise time-step $t$ are projected to the same dimensional latent space by separate feed-forward networks. These features, summed with a sinusoidal positional embedding, are fed into a Transformer encoder-only backbone [51]. The backbone output is projected back to the original motion dimensions via a GCN decoder. Our model is trained both conditionally and unconditionally on text, by randomly masking 10% of the text embeddings. This approach balances diversity and text-fidelity during sampling.
Our method uses the building blocks of MDM [50], but with three key differences: (1) a denoising model that includes variance learning to increase log-likelihood and perform uncertainty estimates, (2) the GCN encoder with learnable graph connectivity, and (3) a learning framework that incorporates contextuality by synchronizing skeletal inputs with initial textual inputs.
@@ 3.1 Problem Formulation
A motion sample can be represented by a temporal skeleton sequence $X=\{p^{i}\}_{i=1}^{N}$ of length $N$ where a frame $p_{i}$ denotes a pose that can be modeled using different joint feature representations depending on the dataset (see Section 4.1). The simplest form that any representation can easily revert to without any loss of information is the joints’ position in 3D space where $p_{i}=\{x(1)_{i},...,x(J)_{i}\}$ with joints $x(j)_{i}\in\mathbb{R}^{M=3}$ and $J$ being the total number of joints. Some parameterizations use rotation matrices ($M=9$), angle-axis ($M=4$), or quaternion ($M=4$) to represent each joint, some also include information such as angular and/or linear velocity.
@@ 3.2 The Variational Diffusion Process
Then, during both training and inference, we use MDMP (see Fig 3) as the decoder—conditioned at each step by the previously mentioned inputs $Y$ and $c$—to progressively denoise $X_{t}$ from a standard Gaussian. Instead of predicting the noise $\epsilon_{0}$ as formulated in DDPM [17], we follow [45] and  [50] and predict the signal itself along with its variance: $\hat{X}_{0},V_{0}=\text{MDMP}(X_{t},t,Y,c)$
@@
One subtlety is that $L_{\text{simple}}$ provides no learning signal for variances, as Ho et al. [17] chose to fix the variance rather than learn it. However, in our framework, we leverage learned variances to generate presence zones with varying confidence levels by reintroducing the variational lower bound term $L_{\mathrm{VLB}}$ as defined in the next subsection.
@@ 3.3 Learning the Variances of the Denoising process
To learn the reverse process variances, our model outputs a vector $V_{0}$ of the same shape as $\hat{X}_{0}$, and-following Nichol et al. [40]—we parameterize the variance as an interpolation between $\beta_{t}$ and $\tilde{\beta}_{t}$ in the log domain by turning this output $V_{0}$ into $\Sigma_{\theta,t}$ as follows:
@@
3.4 Encoding the joint features with GCNs
To encode the spatial pose features, we leverage GCNs [23]. Instead of relying on a predefined sparse graph, we follow Mao et al. [37] and learn the graph connectivity during training, thus essentially learning the dependencies between the different joint trajectories. To this end, we use a fully-connected graph with $N$ nodes, $N$ being the length of the predicted sequence. The strength of the edges in this graph is represented by the weighted adjacency matrix $A\in\mathbb{R}^{N\times N}$. The graph convolutional encoder/decoder then takes as input a matrix $H^{(\text{in})}\in\mathbb{R}^{N\times F}$, where in our case $F=J\times M$ is the number of body joint features. Given the input a matrix $H^{(\text{in})}$, the adjacency matrix $A$ and a set of trainable weights $W\in\mathbb{R}^{F\times\hat{F}}$, a graph convolutional layer outputs a matrix of the form: $H^{(\text{out})}=AH^{(\text{in})}W$. All operations are differentiable with respect to both the adjacency matrix $A$ and the weight matrix $W$, which allows training on both.
@@ 3.5 Predicting Uncertainty
  • Predicted Variance: The final approach uses the learned variance of the predicted distribution of each motion sequence $\Sigma_{\theta,0}$ as the uncertainty factor.
Both the second and third methods produce outputs in the same dimensions as the model, including predictions for root height, root angular and linear velocity, etc. To calculate a single uncertainty index for each joint at each timestep, we average all features associated to the same joint.
@@ 4 Experiments and Results
In this section, we present the experimental setup and evaluation of our proposed model. We describe the dataset used for training and testing, outlines and explains the choice of metrics used for accuracy and uncertainty, and provide details on our model’s implementation. Our comprehensive quantitative and qualitative evaluation, includes comparison with state-of-the-art Text2Motion baselines that propose Motion-Editing re-implemented for a fair comparison with similar conditioning, analysis of uncertainty parameters, and an ablation study to assess the effects of motion-text fusion and architectural design choices.
@@ 4.1 Dataset
To train and evaluate our model, we use the HumanML3D [14] dataset, which is the largest and most diverse collection of scripted human motions. It combines motion sequences from the HumanAct12 [13] and AMASS [36] datasets, processed to standardize the motions to 20 FPS with a maximum length of 10 seconds per sequence. HumanML3D comprises 14,616 motions with 44,970 descriptions, covering 5,371 distinct words, totaling 28.59 hours of motion data with an average length of 7.1s and three textual descriptions per sequence. The dataset is split for training and evaluation. For evaluation, we filter the set to include only motions longer than 3s, allowing us to condition the models on 2.5s of motion and predict at least 0.5s into the future up until more than 5s for longer recorded motions. After filtering, the evaluation set contains 4,328 out of the initial 4,646 motion sequences.
Figure 4: (Left) Temporal evolution of error in predictions. Quantitative Results on HumanML3D over MPJPE [mm]. (Right) 3D Plots of Motion Predictions (orange) vs Ground truth (blue). Motion Sequence example associated to textual prompt:“from a standing position, the person slowly walks in circle, clockwise, then stops”. Paler shades represents earlier frames.
@@ 4.2 Metrics for Accuracy and Uncertainty
To further validate our method we have also added some more metrics in the C.4 Section of the Appendix. First of all, we have re-trained MDM [50] with skeleton data as an input for a direct comparison, demonstrating the efficacy of our architecture. Secondly, one issue with the MPJPE is that it is biased towards one ”ground-truth sequence” and thus heavily penalizes frequency or phase shifts common in longer-term predictions, leading to misleadingly large errors even if motions remain qualitatively realistic. Hence we compared our method with baselines on the NPSS [11] metric which measures similarity in frequency spectra rather than absolute frame-by-frame error, making it better suited to assess the quality of long-term predictions by capturing perceptually relevant motion coherence. Finally, we report results using metrics proposed by Guo et al. [14], such as Frechet Inception Distance (FID), R-Precision, and Multimodality. However, these metrics primarily assess motion quality, semantic alignment with textual input, and variability rather than precise spatial accuracy. Additionally, they depend on pretrained feature extractors not tailored to motion-conditioned predictions.
@@
Figure 5: (Left) Sparsification Error Plot. Quantitative Results of the uncertainty parameters: The Mode Divergence index closely follows the Oracle curve, indicating the strongest alignment between uncertainty estimates and true errors. (Right) Joint Position Evolution over the Denoising Process. During one single diffusion process of specific body joints, the position is progressively denoised until it converges to its final prediction. The fluctuations are used as a parameter for uncertainty.
@@ 4.3 Implementation Details
Our models were trained on an NVIDIA Titan V GPU over 1.7 days and on NVIDIA Tesla V100 GPU over 1.2 days with a batch size of 64. We used 8 layers of the Transformer Encoder with 4 multi-head attention for each, separated by a GeLU activation function and a dropout value of $0.1$. The GCN layer encodes the joint features from $X_{t}$ into a latent dimension of 1024 when learning variances and 512 without learning variances. 1024 corresponds to the concatenation of the joint features of $\hat{X}_{0}$ [512] and $V_{0}$ [512]. To encode the text, we use a frozen CLIP-ViT-B/32 model. Each model was trained for 600K steps, after which a checkpoint was chosen that minimizes the MPJPE metric to be reported. Our generative process is conditioned by a motion input sequence of 50 frames which represents 2.5 seconds at 20FPS. We also set $\lambda=0.001$ to prevent $L_{\text{VLB}}$ from overwhelming $L_{\text{simple}}$. We evaluate our models with guidance-scale $\mu=2.5$ but as discussed in the Motion & Text ablation study Section 4.4 this can be adapted for specific applications (eg. short/long-term predictions).
To evaluate the effectiveness of our multimodal fusion approach, we compare against state-of-the-art Motion Editing baselines MoMask [15], MotionGPT [20] and MDM [50] which are all trained on HumanML3D [14]. We implement their pretrained versions (open-sourced) and compare on the entire test set of HumanML3D using MPJPE. Conversely to Motion Editing, to ensure a fair comparison setting we conditioned each baseline with only the same motion prequel sequence of 50 frames and compared the rest of the predicted sequence to the ground-truth.
4.4 Quantitative & Qualitative Results
Model Accuracy Evaluation over MPJPE: Unlike the implemented baselines MoMask [15], MotionGPT [20] and MDM [50] that treat motion data as a masked input during sampling, our model is trained to leverage it as an additional supervision signal, which we find to lead to significantly enhances performance, especially over longer sequences. In Fig. 4 the temporal evolution chart shows that our model outperforms these baselines in accuracy, with consistently lower MPJPE values over time and a more gradual increase in error. These results are also demonstrated qualitatively in the 3D plots Fig. 4 (Right) (see Appendix & Video for more examples) where our predictions align more closely with the ground truth, especially towards the end of the sequence. Indeed, both baselines’ outputs fail to follow the indicated trajectory (projection of the root joint in the XZ-plane) whereas our model follows the ”circle”, almost aligning with the ground truth on the last frame.